# Entropy Controllable Direct Preference Optimization

### v1

@@ 2 Related Work
When minimizing a certain divergence to bring two probability distributions closer, attention is often given to whether the fitting or divergence is mode-seeking or mode-covering (mass-covering) (Huszár, 2015; Shannon et al., 2020; Ke et al., 2021; Li & Farnia, 2023; Wang et al., 2024a). When fitting a distribution to a multimodal distribution, if the fitted distribution captures one of the modes, this fitting is called mode-seeking. If it covers all the modes, it is termed mode-covering. Accordingly, divergences facilitating such fittings when minimized are similarly referred to as mode-seeking and mode-covering divergences, respectively. The reverse KL divergence, which is used in RLHF and DPO training, is considered mode-seeking compared to forward KL and other $f$-divergence Shannon et al. (2020); Li & Farnia (2023). Policy learning using mode-seeking divergence often performs better than mode-covering divergence (Ke et al., 2021; Wang et al., 2024a). In this study, we propose a new regularizer to replace the minimization of reverse KL divergence in the objective function of DPO, aiming to achieve better performance through enhanced mode-seeking property.
@@ 3.1 Reinforcement Learning from Human Feedbacks (RLHF)
$\begin{split}L(r_{\phi})=-\mathbb{E}_{(x,y_{w},y_{l}\sim\mathcal{D})}[\log\sigma(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l}))].\end{split}$ (2)
@@
$\begin{split}J(\pi_{\theta})=&\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}}\left[r_{\phi}(x,y)\right]-\beta D_{\text{KL}}(\pi_{\theta}(y\mid x)||\pi_{\text{ref}}(y\mid x)),\\ \end{split}$ (3)
@@ 3.2 Directed Preference Optimization (DPO)
$\begin{split}p^{*}(y_{1}\succ y_{2}\mid x)=\sigma\left(\beta\log\frac{\pi^{*}(y_{1}\mid x)}{\pi_{\text{ref}}(y_{1}\mid x)}-\beta\log\frac{\pi^{*}(y_{2}\mid x)}{\pi_{\text{ref}}(y_{2}\mid x)}\right).\end{split}$ (6)
@@
$L_{\text{DPO}}=-\mathbb{E}_{x,y_{w},y_{l}\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\beta\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right].$ (7)
@@ 4.2 H-DPO
$\mathcal{L}(\pi,\lambda,C)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi}[r(x,y)-\alpha\beta\log\pi(y\mid x)+\beta\log\pi_{\text{ref}}(y\mid x)]-\lambda\left(\sum_{y}\pi(y\mid x)-1\right)-\sum_{y}C(y)\pi(y\mid x),$ (10)
@@
$r^{*}(x,y)=\alpha\beta\log\pi^{*}(y\mid x)-\beta\log\pi_{\text{ref}}(y\mid x)+\alpha\beta\log Z(x).$ (12)
@@
$L_{\text{H-DPO}}=-\mathbb{E}_{x,y_{w},y_{l}\sim\mathcal{D}}\left[\log\sigma\left(\alpha\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}-\beta\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right].$ (13)
@@ 5.1 Experimental Setup
We conducted DPO training based on Zephyr-7B-Beta (Tunstall et al., 2023; ). We started from zephyr-7b-sft-full, which is based on Mistral 7B (Jiang et al., 2023) and fine-tuned with UltraChat (Ding et al., 2023). We performed DPO training on it with UltraFeedback (Cui et al., 2023). We evaluated the performance when H-DPO was used instead of standard DPO. The hyperparameters during training were the same as those of Zephyr-7B-beta, except for the variable $\alpha$. The $\alpha$ was varied in the range from 0.8 to 1.2.
The evaluation tasks included diverse grade school math word problems (GSM8K (Cobbe et al., 2021)), coding task (HumanEval (Chen et al., 2021)), multiple-choice question task (MMLU-Pro (Wang et al., 2024d)) and instruction-following task (IFEval (Zhou et al., 2023)). The training was conducted with three different seeds. Further experimental details are provided in Section A.4 and A.3.
@@ 5.2 Performance and Diversity
[…]
DPO ($\alpha=1$) 26.40 ±1.76 28.77 ±0.45 31.83 ±0.17 59.63 ±0.72
[…]
H-DPO ($\alpha=0.9$) 28.83 ±2.32 29.63 ±0.45 32.30 ±0.17 60.93 ±0.50
H-DPO ($\alpha=0.8$) 28.66 ±1.23 27.77 ±0.67 31.93 ±0.19 59.90 ±0.59
[…]
H-DPO ($\alpha=1.2$) 1.718 0.252 0.313 0.690
H-DPO ($\alpha=1.1$) 1.483 0.293 0.296 0.652
DPO ($\alpha=1$) 1.323 0.326 0.289 0.633
H-DPO ($\alpha=0.95$) 1.223 0.339 0.277 0.611
H-DPO ($\alpha=0.9$) 1.113 0.364 0.272 0.590
H-DPO ($\alpha=0.8$) 0.977 0.391 0.268 0.574
5.3 Coverage Evaluation
Figure 4 presents the evaluation results of pass@$k$ for various values of $k$ on the HumanEval benchmark. On HumanEval, there is a negligible difference between models with a small $\alpha$ and standard DPO when $k$ is large. However, interestingly, when $k$ exceeds 100, the results improve for larger $\alpha$ values ($\alpha=1.1$). This suggests that in the evaluation of coverage on HumanEval, the diversity of responses is more crucial than their quality. In standard DPO, as $k$ increases, the best coverage is shown with a temperature of 1. When the temperature exceeds 1, which is higher than the temperature during training, performance drops significantly, indicating that the maximum effective temperature for increasing diversity is 1. As shown in Table 2, increasing $\alpha$ enhances diversity at a temperature of 1. Consequently, for tasks where diversity at a temperature of 1 is crucial, increasing $\alpha$ can improve coverage.
@@ A.1 Preliminary Experiment with Gaussian Distribution
This section details the experiments shown in Figure 1. In this preliminary experiment, we use the proposed regularization $D_{\alpha}(\pi||\pi_{\text{ref}})$, where $D_{\alpha}$ is the same as the Kullback-Leibler divergence $D_{\text{KL}}$ when $\alpha=1$, to estimate the Gaussian distribution $\pi$ that is closest to the Gaussian Mixture Model (GMM) $\pi_{\text{ref}}$. The GMM $\pi_{\text{ref}}$ consists of two components, both weighted equally at 0.5. Each component has a standard deviation of $\sigma_{1}=1$ and $\sigma_{2}=0.8$, respectively, and the experiments were conducted by varying the distance between their means $\mu_{1}$ and $\mu_{2}$ to be 4, 5, and 6. The parameters of the normal distribution $\pi$ were estimated through numerical experiments. Specifically, by minimally altering the location and scale parameters of the normal distribution $\pi$ and using samples from $\pi$, we computed $D_{\alpha}$. The Gaussian distribution for which $D_{\alpha}$ is minimized is denoted as $\hat{\pi}$. The results are plotted in Figure 7, 8 and 9.
Figure 7 represents the case where the means of the GMM components are relatively far apart, making mode-seeking fitting straightforward. In this scenario, using $\alpha=1$, i.e., $D_{\text{KL}}$, also achieves mode-seeking fitting. As $\alpha$ is decreased, the variance of the estimated normal distribution $\hat{\pi}$ decreases gradually. Figure 8 shows the scenario where the distance between the means of the GMM components is reduced compared to Figure 7. Here, $D_{\text{KL}}$ fails to achieve mode-seeking estimation. However, mode-seeking estimation becomes possible when $\alpha$ is slightly decreased, such as to 0.8. In Figure 9, where the mean distance is further reduced, an even smaller $\alpha$, such as 0.6, is required.
These experiments suggest that adjusting $\alpha$ according to the distance between modes makes mode-seeking estimation achievable. Moreover, further decreasing $\alpha$ results in an increasingly smaller variance of the estimated distribution.
@@ A.2 Comparison with $\beta$ tuning
A.3 Other Details
Figure 5: Accuracy on IFEval with various temperatures.
A.4 Evaluation of Diversity
MMLU-Pro was evaluated using the official implementation from Wang et al. (2024d). IFEval and GSM8K were implemented using Gao et al. (2024), where IFEval was evaluated in a 0-shot setting, and GSM8K was evaluated in an 8-shot setting. HumanEval was evaluated using the official implementation from Chen et al. (2021). MMLU-Pro and IFEval were evaluated using one sampling for all test data, and the average accuracy and standard error at a temperature of 0 are shown in Table 1. For GSM8K, 200 test data were used, and for HumanEval, all test data were used, generating 200 responses for each prompt to calculate pass@$k$ based on Chen et al. (2021), as shown in Figure 3 and 4. The results for pass@1 at a temperature of 0.1 are shown in Table 1. The results for varying temperatures in MMLU-Pro, GSM8K, HumanEval, and IFEval are shown in Figure 2, 3, 4 and 5.
Table 3: Average scores of DPO and H-DPO on various tasks.
[…]
@@ A.4 Evaluation of Diversity
(a) $\alpha=0$
(b) $\alpha=0.2$
(c) $\alpha=0.4$
(d) $\alpha=0.6$
(e) $\alpha=0.8$
(f) $\alpha=1$
Figure 7: In the case where $\bm{\mu_{2}-\mu_{1}=6}$, the GMM $\pi_{\text{ref}}$ and $\hat{\pi}$ that minimizes $D_{\alpha}(\pi||\pi_{\text{ref}})$.
(a) $\alpha=0$
(b) $\alpha=0.2$
(c) $\alpha=0.4$
(d) $\alpha=0.6$
(e) $\alpha=0.8$
(f) $\alpha=1$
Figure 8: In the case where $\bm{\mu_{2}-\mu_{1}=5}$, the GMM $\pi_{\text{ref}}$ and $\hat{\pi}$ that minimizes $D_{\alpha}(\pi||\pi_{\text{ref}})$.
(a) $\alpha=0$
(b) $\alpha=0.2$
(c) $\alpha=0.4$
(d) $\alpha=0.6$
(e) $\alpha=0.8$
(f) $\alpha=1$
Figure 9: In the case where $\bm{\mu_{2}-\mu_{1}=4}$, the GMM $\pi_{\text{ref}}$ and $\hat{\pi}$ that minimizes $D_{\alpha}(\pi||\pi_{\text{ref}})$.

### v2

@@ 2 Related Work
When minimizing a certain divergence to bring two probability distributions closer, attention is often given to whether the fitting or divergence is mode-seeking or mode-covering (mass-covering) (Huszár, 2015; Shannon et al., 2020; Ke et al., 2021; Li & Farnia, 2023; Wang et al., 2024a). When fitting a distribution to a multimodal distribution, if the fitted distribution captures one of the modes, this fitting is called mode-seeking. If it covers all the modes, it is termed mode-covering. Accordingly, divergences facilitating such fittings when minimized are similarly referred to as mode-seeking and mode-covering divergences, respectively. The reverse KL divergence, which is used in RLHF and DPO training, is considered mode-seeking compared to forward KL and other $f$-divergence (Shannon et al., 2020; Li & Farnia, 2023). Policy learning using mode-seeking divergence often performs better than mode-covering divergence (Ke et al., 2021; Wang et al., 2024a). In this study, we propose a new regularizer to replace the minimization of reverse KL divergence in the objective function of DPO, aiming to achieve better performance through enhanced mode-seeking property.
@@ 3.1 Reinforcement Learning from Human Feedbacks (RLHF)
$\begin{split}L(r_{\phi})=-\mathbb{E}_{(x,y_{w},y_{l}\sim\mathcal{D})}[\log\sigma(r_{\phi}(x,y_{w})\\ -r_{\phi}(x,y_{l}))].\end{split}$ (2)
@@
$\begin{split}J(\pi_{\theta})=&\ \mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}}\left[r_{\phi}(x,y)\right]\\ &-\beta D_{\text{KL}}(\pi_{\theta}(y\mid x)||\pi_{\text{ref}}(y\mid x)),\\ \end{split}$ (3)
@@ 3.2 Directed Preference Optimization (DPO)
[…]
@@
$\begin{split}L_{\text{DPO}}=-\mathbb{E}_{x,\,y_{w},\,y_{l}\sim\mathcal{D}}\Biggl{[}\log\sigma\biggl{(}&\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{l}\mid x)}\\ \;-\;&\beta\log\frac{\pi_{\text{ref}}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\biggr{)}\Biggr{]}\\[4.0pt] \end{split}$ (7)
@@
[…]
DPO ($\alpha=1$) 26.40 ±1.76 28.77 ±0.45 31.83 ±0.17 59.63 ±0.72
[…]
H-DPO ($\alpha=0.9$) 28.83 ±2.32 29.63 ±0.45 32.30 ±0.17 60.93 ±0.50
H-DPO ($\alpha=0.8$) 28.66 ±1.23 27.77 ±0.67 31.93 ±0.19 59.90 ±0.59
[…]
H-DPO ($\alpha=1.2$) 1.718 0.252 0.313 0.690
H-DPO ($\alpha=1.1$) 1.483 0.293 0.296 0.652
DPO ($\alpha=1$) 1.323 0.326 0.289 0.633
H-DPO ($\alpha=0.95$) 1.223 0.339 0.277 0.611
H-DPO ($\alpha=0.9$) 1.113 0.364 0.272 0.590
H-DPO ($\alpha=0.8$) 0.977 0.391 0.268 0.574
@@ 4.2 H-DPO
$\begin{split}\mathcal{L}(\pi,\lambda,C)=\,&\mathbb{E}_{x\sim\mathcal{D},\,y\sim\pi}\big{[}r(x,y)-\alpha\beta\log\pi(y\mid x)\\ &\qquad\qquad\quad+\beta\log\pi_{\text{ref}}(y\mid x)\big{]}\\ &-\lambda\left(\sum_{y}\pi(y\mid x)-1\right)\\ &-\sum_{y}C(y)\pi(y\mid x)\end{split}$ (10)
@@
$\begin{split}r^{*}(x,y)=\,&\alpha\beta\log\pi^{*}(y\mid x)-\beta\log\pi_{\text{ref}}(y\mid x)\\ &+\alpha\beta\log Z(x)\end{split}$ (12)
@@
[…]
@@ 5.1 Experimental Setup
We conducted DPO training based on Zephyr-7B-Beta (Tunstall et al., 2023, ). We started from zephyr-7b-sft-full, which is based on Mistral 7B (Jiang et al., 2023) and fine-tuned with UltraChat (Ding et al., 2023). We performed DPO training on it with UltraFeedback (Cui et al., 2023). We evaluated the performance when H-DPO was used instead of standard DPO. The hyperparameters during training were the same as those of Zephyr-7B-beta, except for the variable $\alpha$. The $\alpha$ was varied in the range from 0.8 to 1.2. Another model, Llama-3.2-1B (Dubey et al., 2024), was also used for the experiments, and the results are detailed in Section A.3.
The evaluation tasks included diverse grade school math word problems (GSM8K (Cobbe et al., 2021)), coding task (HumanEval (Chen et al., 2021)), multiple-choice question task (MMLU-Pro (Wang et al., 2024d)) and instruction-following task (IFEval (Zhou et al., 2023)). The training was conducted with three different seeds. Further experimental details are provided in Sections A.4 and A.5.
@@ 5.2 Performance and Diversity
5.3 Coverage Evaluation
Figure 4 presents the evaluation results of pass@$k$ for various values of $k$ on the HumanEval benchmark. On HumanEval, there is a negligible difference between models with a small $\alpha$ and standard DPO when $k$ is large. However, interestingly, when $k$ exceeds 100, the results improve for larger $\alpha$ values ($\alpha=1.1$).
5.4 Discussion
In the evaluation of the HumanEval coding task and GSM8K mathematical task, we observed that the optimal values of $\alpha$ differed between these two task categories. This discrepancy can be attributed to differences in task characteristics, which necessitate distinct sampling temperatures for effective generation. In mathematical tasks, where there is a single correct answer and precise reasoning is required, more deterministic sampling with a lower temperature is preferable. In these cases, values of $\alpha$ less than 1 are suited, facilitating more precise generations. Conversely, in coding tasks, multiple valid answers typically exist, and generating diverse outputs increases the likelihood of producing correct responses. As a result, a sampling temperature of 1 is more suitable for pass@$k$ evaluations in such scenarios. Note that when the temperature exceeds the training value of 1, a significant decline in performance is observed. In such cases, values of $\alpha$ greater than 1 further enhance diversity, as shown in Table 2, improving the probability of generating correct responses in pass@$k$ evaluations.
In summary, for tasks requiring accuracy and utilizing a temperature lower than 1, an $\alpha$ value slightly less than 1, such as 0.9 or 0.95, is appropriate. Conversely, for tasks emphasizing diversity and employing a temperature of 1, using an $\alpha$ value greater than 1, such as 1.1, yields better results.
As suggested by Figures 3 and 4, a practical approach to tuning the $\alpha$ parameter is to first train the model using the standard DPO setting ($\alpha=1$) and then evaluate the performance changes by varying the temperature. For HumanEval with smaller $k$ values and GSM8K, performance improves when the temperature is slightly reduced from 1, indicating that more accurate outputs are preferable, and this improvement aligns with lowering $\alpha$. Conversely, for HumanEval with larger $k$ values, performance degrades as the temperature decreases from 1, suggesting that diversity is critical in such cases, which explains the relatively better performance with $\alpha>1$. In this way, $\alpha$ tuning can be guided by observing whether performance improves or declines when the temperature deviates from 1.
@@ 6 Conclusion
Impact Statement
As this paper primarily focuses on the algorithmic contributions to fine-tuning language models using DPO, its direct societal impact is limited. However, the application of our methodology, particularly in the context of RLHF, requires careful consideration of the feedback process. The individuals providing feedback play a crucial role in shaping the behavior of the language model. Ensuring that the feedback discourages harmful, malicious, or unethical outputs is essential for aligning the model with societal norms and ethical standards.
@@ A.1 Preliminary Experiment with Gaussian Distribution
This section details the experiments shown in Figure 1. In this preliminary experiment, we use the proposed regularization $D_{\alpha}(\pi||\pi_{\text{ref}})$, where $D_{\alpha}$ is the same as the Kullback-Leibler divergence $D_{\text{KL}}$ when $\alpha=1$, to estimate the Gaussian distribution $\pi$ that is closest to the Gaussian Mixture Model (GMM) $\pi_{\text{ref}}$. The experiments were conducted with GMMs $\pi_{\text{ref}}$ consisting of 2, 3, and 4 Gaussian components, and the results are shown in Figures 5, 7 and 8, respectively. For any $\pi_{\text{ref}}$, the weights of the components are equal, and the standard deviations are 1 and 0.8 for the case of two components, 1, 0.8, and 0.5 for the case of three components, and 1, 0.8, 0.5, and 0.3 for the case of four components. In those figures, the results of varying the interval between the means of the Gaussian components are displayed in separate rows.
In the upper row of Figure 5, we observe that when $\alpha=1$, i.e. using the KL divergence, the fitting becomes mode-covering. When $\alpha$ is reduced to 0.6, it successfully achieves mode-seeking fitting. In the middle row, where the interval between the means of the components is larger, making mode-seeking fitting more feasible, mode-seeking fitting is observed at $\alpha=0.8$. In the bottom row, where the interval is even larger, mode-seeking fitting occurs even when minimizing the KL divergence, although the fitting targets the Gaussian with the larger variance on the left. As $\alpha$ decreases, the fitting shifts to the Gaussian on the right, which has smaller variance and higher probability.
In Figure 6, the values of $D_{\alpha}(\pi||\pi_{\text{ref}})$ are represented using color as the location and scale parameters of the Gaussian distribution $\hat{\pi}$ are varied. As $\alpha$ decreases, the $D_{\alpha}(\pi||\pi_{\text{ref}})$ values for mode-seeking $\hat{\pi}$ become smaller compared to those for mode-covering $\hat{\pi}$.
Similar results are observed in Figures 7 and 8 for cases with 3 and 4 Gaussian components. When minimizing the KL divergence, the fitting tends to be mode-covering or targets the component with larger variance. However, reducing $\alpha$ results in the fitting successfully targeting the region with the highest probability in all cases. As $\alpha$ decreases further, the variance of $\hat{\pi}$ also becomes smaller.
(a) $\alpha=0$
(b) $\alpha=0.2$
(c) $\alpha=0.4$
(d) $\alpha=0.6$
(e) $\alpha=0.8$
(f) $\alpha=1$
Figure 5: $\pi_{\text{ref}}$ is a GMM composed of two normal distributions, and $\hat{\pi}$ represents the normal distribution that minimizes $D_{\alpha}(\pi||\pi_{\text{ref}})$. The upper, middle, and bottom rows correspond to cases where the mean intervals between components are 4, 5, and 6, respectively. The standard deviations of each component are 1 and 0.8 from left to right.
Figure 6: Values of $D_{\alpha}(\pi||\pi_{\text{ref}})$ for the normal distribution $\pi$ with various location and scale parameters in the experiment shown in Figure 5. For visibility, $\min(3,\ln D_{\alpha}(\pi||\pi_{\text{ref}})-\ln D_{\alpha}(\hat{\pi}||\pi_{\text{ref}}))$ is plotted. The red star indicates the parameters of $\hat{\pi}$ that minimize $D_{\alpha}(\pi||\pi_{\text{ref}})$, and these values are used to plot Figure 5.
(a) $\alpha=0$
(b) $\alpha=0.2$
(c) $\alpha=0.4$
(d) $\alpha=0.6$
(e) $\alpha=0.8$
(f) $\alpha=1$
Figure 7: $\pi_{\text{ref}}$ is a GMM composed of two normal distributions, and $\hat{\pi}$ represents the normal distribution that minimizes $D_{\alpha}(\pi||\pi_{\text{ref}})$. The upper, middle, and bottom rows correspond to cases where the mean intervals between components are 3, 5, and 7, respectively. The standard deviations of each component are 1, 0.8 and 0.5 from left to right.
(a) $\alpha=0$
(b) $\alpha=0.2$
(c) $\alpha=0.4$
(d) $\alpha=0.6$
(e) $\alpha=0.8$
(f) $\alpha=1$
Figure 8: $\pi_{\text{ref}}$ is a GMM composed of two normal distributions, and $\hat{\pi}$ represents the normal distribution that minimizes $D_{\alpha}(\pi||\pi_{\text{ref}})$. The upper, middle, and bottom rows correspond to cases where the mean intervals between components are 3, 5, and 7, respectively. The standard deviations of each component are 1, 0.8, 0.5 and 0.3 from left to right.
@@ A.2 Comparison with $\beta$ tuning
A.3 Experiments with Llama-3.2-1B
To further demonstrate the applicability of H-DPO across different models, we conducted experiments using Llama-3.2-1B (Dubey et al., 2024). To differentiate these experiments from those performed with Zephyr, we utilized a different dataset, the Anthropic HH dataset (Bai et al., 2022). The experimental setup was consistent with Rafailov et al. (2023), where Llama-3.2-1B underwent SFT using only the preference completions from the dataset, followed by fine-tuning with H-DPO. The value of $\beta$ was set to 0.01, while all other hyperparameters matched those used in Rafailov et al. (2023).
The results of the experiments conducted on four tasks are presented in Table 3. Given the difficulty of the tasks and the inherently low performance of the base model, consistent improvements were not observed in HumanEval. However, we did observe performance improvements in other tasks.
Table 3: Average scores of DPO and H-DPO with different $\alpha$ values on various tasks when using the Llama-3.2-1B model.
[…]
DPO ($\alpha=1$) 4.97 ±0.31 2.73 ±1.43 14.20 ±0.14 22.60 ±0.08
[…]
H-DPO ($\alpha=0.9$) 4.40 ±0.19 0.57 ±0.28 14.13 ±0.12 23.67 ±0.10
A.4 Evaluation of Diversity
Figure 10: Accuracy on IFEval with various temperatures.
A.5 Other Details
MMLU-Pro was evaluated using the official implementation from Wang et al. (2024d). IFEval and GSM8K were implemented using Gao et al. (2024), where IFEval was evaluated in a 0-shot setting, and GSM8K was evaluated in an 8-shot setting. HumanEval was evaluated using the official implementation from Chen et al. (2021). MMLU-Pro and IFEval were evaluated using one sampling for all test data, and the average accuracy and standard error at a temperature of 0 are shown in Tables 1 and 3. For GSM8K, 200 test data were used, and for HumanEval, all test data were used, generating 200 responses for each prompt to calculate pass@$k$ based on Chen et al. (2021), as shown in Figures 3 and 4. The results for pass@1 at a temperature of 0.1 are shown in Tables 1 and 3. The results for varying temperatures in MMLU-Pro, GSM8K, HumanEval, and IFEval are shown in Figures 2, 3, 4 and 10.
Table 4: Average scores of DPO and H-DPO on various tasks.
[…]